OpenStudy (anonymous):

35y^2-875/49y^2-1225 simplify by removing factors of 1

OpenStudy (anonymous):

35(y^2-25)/49(y^2-25) = 35/49 = 5/7

OpenStudy (anonymous):

yes right factor it